# Fe contribution to the magnetic anisotropy of $L{\rm 1_{0}}$-ordered FePt thin films studied by angle-dependent x-ray magnetic circular dichroism

### v1

@@ Abstract
Magnetic thin films with perpendicular magnetic anisotropy (PMA) have been extensively studied since they are essential for applications in high-density magnetic recording media and novel spintronics devices. Dieny and Chshiev (2017); Tudu and Tiwari (2017); Miwa et al. (2018); Chen et al. (2022) Among them, $L1_{0}$-ordered FePt has drawn particularly strong attention because it shows PMA with exceptionally high magnetocrystalline anisotropy (MCA) energy of $\sim 10^{6}$$10^{7}$ J/m3 (Refs. Weller et al., 2000; Seki et al., 2011a, 2017). By adjusting the deposition and post-annealing temperatures, one can control the degree of the long range chemical order ($S$) in a wide range, Seki et al. (2011a); Shima et al. (2002) a useful property for realizing desired device characteristics. Voltage-controlled magnetic anisotropy of the $L1_{0}$-ordered FePt thin films has also been a topical issue in recent years. Weisheit et al. (2007); Seki et al. (2011b); Miwa et al. (2018); Nozaki et al. (2019); Chen et al. (2022); Miwa et al. (2017)
@@
In order to elucidate the origin of the PMA in $L1_{0}$-FePt, it is essential to experimentally probe the orbital asphericity as well as $m_{\text{orb}}$ of the Fe $3d$ and Pt $5d$ electrons. XMCD is an element-specific technique which can separately probe the orbital and spin magnetic moments of ferromagnetic materials by using XMCD sum rules. Thole et al. (1992); Carra et al. (1993) Moreover, it is known that the deduced spin magnetic moment from the XMCD sum rule, hereafter called as ‘effective’ spin magnetic moment ($m_{\text{spin}}^{\text{eff}}$), contains an additional term called ‘magnetic dipole moment’ ($m_{\text{T}}$), which is directly connected to the orbital asphericity of the spin-polarized electrons. Carra et al. (1993); Stöhr and König (1995); Dürr and van der Laan (1996); Stöhr (1999) It is known that the magnetic dipole moment $m_{\text{T}}$ can be separated from the spin magnetic moment ($m_{\text{spin}}$) by measuring the magnetic-field angle dependence of the XMCD spectra. Dürr and van der Laan (1996); Dürr et al. (1997); van der Laan et al. (2010); Shibata et al. (2018); Sakamoto et al. (2021) In the present study, in order to reveal the orbital asphericity of the Fe $3d$ electrons, we have performed angle-dependent XMCD (AD-XMCD) experiments at the Fe $L_{2,3}$ edge of FePt thin films with different degrees of $L1_{0}$ order and hence different MCA energies. From the magnetic field-angle dependence of $m_{\text{spin}}^{\text{eff}}$, it has been shown that the spin-polarized electrons of Fe preferentially occupy orbitals elongated along the $z$ axis (the out-of-plane direction). The deduced MCA energy due to the orbital asphericity is negative, and its magnitude is larger than that due to the OMA. This means that the Fe sites favor the in-plane magnetic easy axis and that Pt should play a dominant role in the PMA of $L1_{0}$-ordered FePt thin films.
@@
Figure 1: Fe $L_{2,3}$-edge x-ray absorption spectroscopy (XAS) and angle-dependent x-ray magnetic circular dichroism (AD-XMCD) measurements. (a) Schematic illustration of the experimental geometry. The angles of the incident x ray ($\theta_{\text{inc}}$), the external magnetic field ($\theta_{\bm{H}}$) and the magnetization ($\theta_{\bm{M}}$) are defined with respect to the sample normal. $\theta_{\text{inc}}$ is fixed at $60^{\circ}$ in the present measurements. $\theta_{\bm{H}}$ can be varied in the range of $-{90}^{\circ}\leq\theta_{\bm{H}}\leq{90}^{\circ}$. Note that the external magnetic field $\bm{H}$ and the magnetization $\bm{M}$ are not always parallel in the presence of magnetic anisotropy. (b–d) XAS and (e–g) AD-XMCD spectra of the FePt thin films with $L1_{0}$ order degrees $S=0$ (b,e), $S=0.3$ (c,f) and $S=0.4$ (d,g). Linear and smoothed two-step backgrounds have been subtracted from the raw XAS spectra. All the XAS spectra have been normalized to the height of the $L_{3}$ peak at $\sim 708\ \text{eV}$. $\theta_{\bm{H}}=60^{\circ}$ corresponds to the geometry where the incident x ray and the magnetic field are parallel [See Panel (a)]. XMCD intensity at the $L_{3}$ edge is proportional to the projected magnetic moment $\hat{\bm{P}}\cdot\bm{M}$ shown in Panel (a).
@@
Table 1: Sample-preparation conditions, the degree of the $L1_{0}$ order $S$, and the magnetocrystalline anisotropy (MCA) energy deduced from magnetization curves $K_{\text{u}}^{M\text{-}H}$. We note that the magnetic anisotropy due to the demagnetizing field (shape anisotropy) has been corrected when calculating $K_{\text{u}}^{M\text{-}H}$.
@@
Deposition temperature $T_{\text{S}}$ (${}^{\circ}\mathrm{C}$) Annealing temperature $T_{A}$ (${}^{\circ}\mathrm{C}$) $L1_{0}$ order degree $S$ $K_{\text{u}}^{M\text{-}H}$ (MJ/m3)
300 350 0.4 1.3
300 0.3 0.7
RT 0 0.3
FePt thin films were grown on MgO (100) substrates by the ultrahigh vacuum magnetron-sputtering method. The details of the sample preparation methods are described in Ref. Seki et al., 2011a. The stacked structure of the sample was MgO substrate/Fe (1 nm)/Au (30 nm)/FePt (20 nm)/Au (2 nm). The Au 2 nm layer at the top was deposited as a capping layer in order to prevent the sample from oxidization. The 30-nm-thick Au layer was deposited as a buffer layer to ensure that the lattice constant of FePt is not influenced by that of the substrate, and the 1-nm-thick Fe layer is the seed layer. One can control the degree of the $L1_{0}$ order $S$ (see Refs. Warren, 1990 and Cebollada et al., 1994 for the definition of $S$) by changing the substrate temperature during deposition $T_{\text{S}}$ or the subsequent annealing temperature after the deposition $T_{\text{A}}$. Seki et al. (2011a) Table 1 summarizes the sample-preparation conditions, the degree of the $L1_{0}$ order $S$ deduced from the intensity of the (001) x-ray diffraction peak, and the MCA energy deduced from magnetization measurements $K_{\text{u}}^{M\text{-}H}$ (see Supplementary Materials for the x-ray diffraction profiles and the magnetization curves).
@@
The Fe $L_{2,3}$-edge x-ray absorption spectroscopy (XAS) and AD-XMCD measurements were performed at helical undulator beamline BL-16A of Photon Factory, High Energy Accelerator Research Organization (KEK-PF). Figure 1(a) is the schematic illustration of the measurement geometry. The x-ray incident angle $\theta_{\text{inc}}$ was fixed at $60^{\circ}$ from the sample normal and the magnetic field direction relative to the sample normal $\theta_{\bm{H}}$ was varied in the range of $-90^{\circ}\leq\theta_{\bm{H}}\leq+90^{\circ}$ (where the sample normal is defined as $0^{\circ}$) using two orthogonal pairs of superconducting magnets. Furuse et al. (2013); Shibata et al. (2018); Sakamoto et al. (2021); Nonaka et al. (2021) From the different angular dependencies between the magnetic dipole moment $m_{\text{T}}$ and the spin magnetic moment $m_{\text{spin}}$, one can deduce $m_{\text{T}}$ by AD-XMCD. Stöhr and König (1995); Dürr and van der Laan (1996); Shibata et al. (2018); Sakamoto et al. (2021) The maximum magnitude of the applied magnetic field was 1 T. All the measurements were performed at room temperature. Absorption signals were collected in the total electron-yield method. The hole number $n_{h}$ in the XMCD sum rule Carra et al. (1993) was assumed to be 3.4 based on a previous band-structure calculation.Schmitz-Antoniak et al. (2011)
@@
Figure 2: Magnetic field angle dependence of effective spin magnetic moment projected on the light axis ($\hat{\bm{P}}\cdot\bm{m}_{\text{spin}}^{\text{eff}}$). (a) Experimental data for the film with $S=0.4$ (red square) and simulated angular dependencies based on the Stoner-Wohlfarth model without magnetocrystalline anisotropy (MCA, blue dotted curve) and with MCA and $\braket{Q_{zz}}$ (black curve). (b) Experimental and simulated angular dependencies of $\hat{\bm{P}}\cdot\bm{m}_{\text{spin}}^{\text{eff}}$ for various $L1_{0}$ order degrees $S$. Both $K_{\text{u}}$ and $\braket{Q_{zz}}$ are incorporated in the simulation.
@@
where $M$ is the total magnetization (the sum of the spin and orbital magnetic moments of Fe and Pt). The three terms in Eq. (1) represent the Zeeman energy, the magnetic shape anisotropy of the film, and the MCA energy. By minimizing $E_{m}$ with respect to $\theta_{\bm{M}}$, one can deduce the value of $\theta_{\bm{M}}$ for each angle $\theta_{\bm{H}}$, and thus the effective spin magnetic moment can be deduced by the following equation: Shibata et al. (2018); Sakamoto et al. (2021); Nonaka et al. (2021)
@@
where $\theta_{\text{inc}}=60^{\circ}$ is the x-ray incidence angle. In the simulation, we have assumed that the angular dependence of $m_{\text{spin}}+m_{\text{orb}}$ is negligible and that the ratio of $m_{\text{spin}}^{\text{Pt}}$ to $m_{\text{spin}}^{\text{Fe}}$ is constant ($m_{\text{spin}}^{\text{Pt}}/m_{\text{spin}}^{\text{Fe}}\simeq 0.11$)Ikeda et al. (2017). The curves in Fig. 2(a) represent thus simulated angular dependencies with and without MCA energy ($K_{\text{u}}$) and quadrupolar moment $\braket{Q_{zz}}$. One can see that the simulation with both $K_{\text{u}}$ and $\braket{Q_{zz}}$ reasonably reproduces the experimental data. Figure 2(b) shows the experimental and simulated angular dependencies of the effective spin magnetic moment for all the samples based on simulation. The fitted values of $m_{\text{spin}}$, $K_{\text{u}}$, and $\braket{Q_{zz}}$ are summarized in Table 2. The MCA energy $K_{\text{u}}$ increases with $S$, consistent with the magnetization measurements (Fig. S2 of Supplementary Materials) and the previous magnetization studies Okamoto et al. (2002); Seki et al. (2011a). One can see that $\braket{Q_{zz}}$ is negative for the films which exhibit PMA ($S=0.3,0.4$) and that the magnitude of $|\braket{Q_{zz}}|$ increases with $S$ and $K_{\text{u}}$. Figure 3 summarizes the $S$ dependence of $\braket{Q_{zz}}$. The negative $\braket{Q_{zz}}$ values show that the charge density in Fe ions is elongated along the out-of-plane direction, as shown in the bottom of Fig. 3. We note that the decrease of $m_{\text{spin}}$ with $S$ in Table 2 is consistent with the previous experimental studyIkeda et al. (2017) and the theoretical calculations based on the coherent potential approximation. Staunton et al. (2004)
@@
|$\ket{\rm ex}$ $=$ $c^{\dagger}_{n_{2},\sigma_{2}}c_{n_{1},\sigma_{1}}\ket{\rm gr}\qquad(\sigma_{1},\sigma_{2}=\pm 1/2),$
@@
where $\zeta$ is the SOC constant and $\Delta E_{\text{ex}}$ is the exchange splitting between the majority- and minority-spin bands. The first term, originating from spin-conserved excitations, is proportional to the orbital moment anisotropy (OMA) between the film out-of-plane and in-plane directions. The second term, originating from spin-flip excitations, is proportional to the anisotropic charge distribution $\braket{Q_{zz}}$. Using the OMA deduced in our previous XMCD measurements Ikeda et al. (2017) and the $\braket{Q_{zz}}$ values deduced in the present study (Table 2 and Fig. 3), we have estimated the contributions of Fe to the MCA of FePt thin films and have plotted them in Fig. 4 as functions of the $L1_{0}$ order degree $S$. Here, $\zeta$ and $\Delta E_{\text{ex}}$ of Fe have been assumed to be 0.054 eV and 3 eV, respectively van der Laan (1998); Kota and Sakuma (2012). The OMA is positive and gives a positive contribution to the PMA. In contrast, the negative spin-flip term of Fe favors the in-plane magnetic easy axis. Besides, its magnitude is even larger than that of the spin-conservation term. Therefore, the sum of the spin-conservation term and the spin-flip term is also negative with increasing $S$, as shown in Fig. 4.
Figure 4: Contribution to the MCA energy originating from the orbital moment anisotropy (spin-conservation term, blue) Ikeda et al. (2017), and the charge-density anisotropy (spin-flip term, red). The sum of the two terms (total MCA energy) is also shown (green).
@@
The present results also indicate that the PMA of $L1_{0}$ FePt thin films mainly originates from the Pt sites. Our previous XMCD study at the Pt $L_{2,3}$ edge Ikeda et al. (2017) shows that the OMA of Pt is negative ($m_{\text{orb}}^{\perp}<m_{\text{orb}}^{\parallel}$), suggesting that the OMA favors the in-plane magnetic easy axis. Therefore, it follows that the PMA of the $L1_{0}$ FePt thin films is primarily due to the orbital asphericity of the Pt sites. These results are consistent with the theoretical calculation using the Green’s function method Solovyev, Dederichs, and Mertig (1995). To verify this hypothesis, AD-XMCD experiments at the Pt $L_{2,3}$ edge using hard x rays are desired, which remain to be performed in the future.
[…]

### v2

@@ Abstract
Magnetic thin films with perpendicular magnetic anisotropy (PMA) have been extensively studied since they are essential for applications in high-density magnetic recording media and novel spintronics devices. Dieny and Chshiev (2017); Tudu and Tiwari (2017); Miwa et al. (2019); Chen et al. (2022) Among them, $L1_{0}$-ordered FePt has drawn particularly strong attention because it shows PMA with exceptionally high magnetocrystalline anisotropy (MCA) energy of $\sim 10^{6}$$10^{7}$ J/m3. Weller et al. (2000); Seki et al. (2011a, 2017) By adjusting the deposition and post-annealing temperatures, one can control the degree of the long range chemical order ($S$) in a wide range, Seki et al. (2011a); Shima et al. (2002) a useful property for realizing desired device characteristics. Voltage-controlled magnetic anisotropy of the $L1_{0}$-ordered FePt thin films has also been a topical issue in recent years. Weisheit et al. (2007); Seki et al. (2011b); Miwa et al. (2019); Nozaki et al. (2019); Chen et al. (2022); Miwa et al. (2017)
@@
In order to elucidate the origin of the PMA in $L1_{0}$-FePt, it is essential to experimentally probe the orbital asphericity as well as $m_{\text{orb}}$ of the Fe $3d$ and Pt $5d$ electrons. XMCD is an element-specific technique which can separately probe the orbital and spin magnetic moments of ferromagnetic materials by using XMCD sum rules. Thole et al. (1992); Carra et al. (1993) Moreover, it is known that the deduced spin magnetic moment from the XMCD sum rule, hereafter called as ‘effective’ spin magnetic moment ($m_{\text{spin}}^{\text{eff}}$), contains an additional term called ‘magnetic dipole moment’ ($m_{\text{T}}$), which is directly connected to the orbital asphericity of the spin-polarized electrons. Carra et al. (1993); Stöhr and König (1995); Dürr and van der Laan (1996); Stöhr (1999); Dürr et al. (1997) It is known that the magnetic dipole moment $m_{\text{T}}$ can be separated from the spin magnetic moment ($m_{\text{spin}}$) by measuring the magnetic-field angle dependence of the XMCD spectra. Dürr and van der Laan (1996); Stöhr (1999); Dürr et al. (1997); van der Laan et al. (2010); Shibata et al. (2018); Sakamoto et al. (2021) In the present study, in order to reveal the orbital asphericity of the Fe $3d$ electrons, we have performed angle-dependent XMCD (AD-XMCD) experiments at the Fe $L_{2,3}$ edge of FePt thin films with different degrees of $L1_{0}$ order and hence different MCA energies. From the magnetic field-angle dependence of $m_{\text{spin}}^{\text{eff}}$, it has been shown that the spin-polarized electrons of Fe preferentially occupy orbitals elongated along the $z$ axis (the out-of-plane direction). The deduced MCA energy due to the orbital asphericity is negative, and its magnitude is larger than that due to the OMA. This means that the Fe sites favor the in-plane magnetic easy axis and that Pt should play a dominant role in the PMA of $L1_{0}$-ordered FePt thin films.
@@
Figure 1: Fe $L_{2,3}$-edge x-ray absorption spectroscopy (XAS) and angle-dependent x-ray magnetic circular dichroism (AD-XMCD) measurements. (a) Schematic illustration of the experimental geometry. The angles of the incident x ray ($\theta_{\text{inc}}$), the external magnetic field ($\theta_{\bm{H}}$) and the magnetization ($\theta_{\bm{M}}$) are defined with respect to the sample normal. $\theta_{\text{inc}}$ is fixed at $60^{\circ}$ in the present measurements. $\theta_{\bm{H}}$ can be varied in the range of ${90}^{\circ}\leq\theta_{\bm{H}}\leq{90}^{\circ}$. Note that the external magnetic field $\bm{H}$ and the magnetization $\bm{M}$ are not always parallel in the presence of magnetic anisotropy. (b–d) XAS and (e–g) AD-XMCD spectra of the FePt thin films with $L1_{0}$ order degrees $S=0$ (b,e), $S=0.3$ (c,f) and $S=0.4$ (d,g). Linear and smoothed two-step backgrounds have been subtracted from the raw XAS spectra. All the XAS spectra have been normalized to the height of the $L_{3}$ peak at $\sim 708\ \text{eV}$. $\theta_{\bm{H}}=60^{\circ}$ corresponds to the geometry where the incident x ray and the magnetic field are parallel [See Panel (a)]. XMCD intensity at the $L_{3}$ edge is proportional to the projected magnetic moment $\hat{\bm{P}}\cdot\bm{M}$ shown in Panel (a).
@@
Table 1: Sample-preparation conditions, the degree of the $L1_{0}$ order $S$, magnetic anisotropy energy due to the demagnetizing field (shape anisotropy) $\frac{\mu_{0}}{2}M_{\text{s}}^{2}$, and the magnetocrystalline anisotropy (MCA) energy deduced from magnetization curves $K_{\text{u}}^{M\text{-}H}$. We note that $K_{\text{u}}^{M\text{-}H}$ is the sum of the shape anisotropy energy and the effective magnetic anisotropy energy $K_{\text{eff}}$ calculated directly from the $M$-$H$ curves.
@@
Deposition temperature $T_{\text{S}}$ ($^{\circ}\mathrm{C}$) Annealing temperature $T_{A}$ ($^{\circ}\mathrm{C}$) $L1_{0}$ order degree $S$ Shape anisotropy energy $\frac{\mu_{0}}{2}M_{\text{s}}^{2}$ $K_{\text{u}}^{M\text{-}H}$ (MJ/m3)
300 350 0.4 0.80 1.3
300 0.3 0.71 0.7
RT 0 0.51 0.3
[…]
@@
The Fe $L_{2,3}$-edge x-ray absorption spectroscopy (XAS) and AD-XMCD measurements were performed at helical undulator beamline BL-16A of Photon Factory, High Energy Accelerator Research Organization (KEK-PF). Figure 1(a) is the schematic illustration of the measurement geometry. The x-ray incident angle $\theta_{\text{inc}}$ was fixed at $60^{\circ}$ from the sample normal and the magnetic field direction relative to the sample normal $\theta_{\bm{H}}$ was varied in the range of $90^{\circ}\leq\theta_{\bm{H}}\leq+90^{\circ}$ (where the sample normal is defined as $0^{\circ}$) using two orthogonal pairs of superconducting magnets. Furuse et al. (2013); Shibata et al. (2018); Sakamoto et al. (2021); Nonaka et al. (2021) From the different angular dependencies between the magnetic dipole moment $m_{\text{T}}$ and the spin magnetic moment $m_{\text{spin}}$, one can deduce $m_{\text{T}}$ by AD-XMCD. Stöhr and König (1995); Dürr and van der Laan (1996); Shibata et al. (2018); Sakamoto et al. (2021) The maximum magnitude of the applied magnetic field was 1 T. In order for the potential experimental artifacts to be cancelled out, the XMCD measurements were performed for the positive and negative (reversed) magnetic fields at each magnetic field angle $\theta_{\bm{H}}$ and the two obtained XMCD spectra were averaged. All the measurements were performed at room temperature. Absorption signals were collected in the total electron-yield method. On the application of the XMCD sum rules, we first subtracted the polygonal line bent at the Fe $L_{3}$ XAS peak and then a two-step white line background composed of arctangent functions. On the application of the spin sum rule Carra et al. (1993), the boundary between the Fe $L_{3}$ and $L_{2}$ edges was set to be $h\nu=718$ eV and the hole number $n_{h}$ was assumed to be 3.4 based on a previous band-structure calculation.Schmitz-Antoniak et al. (2011) Details of the sum rule analysis have been described in one of our previous report. Shibata et al. (2021) We note that the arbitrariness of these procedures on applying the XMCD sum rules can result in systematic errors of $\sim\pm(10-15)\%$ in the deduced magnetic moments. Shibata et al. (2021)
@@
Figure 2: Magnetic field angle dependence of effective spin magnetic moment projected on the light axis ($\hat{\bm{P}}\cdot\bm{m}_{\text{spin}}^{\text{eff}}$). (a) Experimental data for the film with $S=0.4$ (red square) and simulated angular dependencies based on the Stoner-Wohlfarth model without magnetocrystalline anisotropy (MCA) and $\braket{Q_{zz}}$ (blue dotted curve), with MCA but without $\braket{Q_{zz}}$ (green dotted curve), and with both MCA and $\braket{Q_{zz}}$ (black curve). (b) Experimental and simulated angular dependencies of $\hat{\bm{P}}\cdot\bm{m}_{\text{spin}}^{\text{eff}}$ for various $L1_{0}$ order degrees $S$. Both $K_{\text{u}}$ and $\braket{Q_{zz}}$ are incorporated in the simulation.
@@
where $M$ is the total magnetization (the sum of the spin and orbital magnetic moments of Fe and Pt). The three terms in Eq. (1) represent the Zeeman energy, the general magnetic shape anisotropy for thin films, and the MCA energy. By minimizing $E_{m}$ with respect to $\theta_{\bm{M}}$, one can deduce the value of $\theta_{\bm{M}}$ for each angle $\theta_{\bm{H}}$, and thus the effective spin magnetic moment can be deduced by the following equation: Shibata et al. (2018); Sakamoto et al. (2021); Nonaka et al. (2021)
@@
where $\theta_{\text{inc}}=60^{\circ}$ is the x-ray incidence angle. In the simulation, we have assumed that the angular dependence of $m_{\text{spin}}+m_{\text{orb}}$ is negligible and that the ratio of $m_{\text{spin}}^{\text{Pt}}$ to $m_{\text{spin}}^{\text{Fe}}$ is constant ($m_{\text{spin}}^{\text{Pt}}/m_{\text{spin}}^{\text{Fe}}\simeq 0.11$)Ikeda et al. (2017). The curves in Fig. 2(a) represent thus simulated angular dependencies under various assumptions: (i) both MCA ($K_{\text{u}}$) and quadrupolar moment $\braket{Q_{zz}}$ are absent, (ii) MCA is incorporated but $\braket{Q_{zz}}$ is absent, and (iii) both MCA and $\braket{Q_{zz}}$ are taken into account. Simulation (i) clearly deviates from the experimental data, showing that one needs to incorporate MCA in order to reproduce the experimental angular dependence. Comparing simulations (ii) and (iii), the latter (simulation with $\braket{Q_{zz}}$) better agrees with the experimental data. Figure 2(b) shows the experimental and simulated angular dependencies of the effective spin magnetic moment for all the samples based on simulation. All the data could be well reproduced based on the Stoner-Wohlfarth model, suggesting that the assumption of the magnetic single domain is valid. The fitted values of $m_{\text{spin}}$, $K_{\text{u}}$, and $\braket{Q_{zz}}$ are summarized in Table 2. The MCA energy $K_{\text{u}}$ increases with $S$, consistent with the magnetization measurements (Fig. S2 of Supplementary Materials) and the previous magnetization studies Okamoto et al. (2002); Seki et al. (2011a). One can see that $\braket{Q_{zz}}$ is negative for the films which exhibit PMA ($S=0.3,0.4$) and that the magnitude of $|\braket{Q_{zz}}|$ increases with $S$ and $K_{\text{u}}$. Figure 3 summarizes the $S$ dependence of $\braket{Q_{zz}}$. The negative $\braket{Q_{zz}}$ values show that the charge density in Fe ions is elongated along the out-of-plane direction, as shown in the bottom of Fig. 3. We note that the decrease of $m_{\text{spin}}$ with $S$ in Table 2 is consistent with the previous experimental studyIkeda et al. (2017) and the theoretical calculations based on the coherent potential approximation. Staunton et al. (2004)
@@
where $\zeta$ is the SOC constant and $\Delta E_{\text{ex}}$ is the exchange splitting between the majority- and minority-spin bands. The first term, originating from spin-conserved excitations, is proportional to the orbital moment anisotropy (OMA) between the film out-of-plane and in-plane directions. The second term, originating from spin-flip excitations, is proportional to the anisotropic charge distribution $\braket{Q_{zz}}$. Using the OMA and the $\braket{Q_{zz}}$ values deduced in the present study (Table 2 and Fig. 3), we have estimated the contributions of Fe to the MCA of FePt thin films and have plotted them in Fig. 4 as functions of the $L1_{0}$ order degree $S$. Here, $\zeta$ and $\Delta E_{\text{ex}}$ of Fe have been assumed to be 0.054 eV and 3 eV, respectively van der Laan (1998); Kota and Sakuma (2012). The OMA is positive and gives a small positive contribution to the PMA. In contrast, the negative spin-flip term of Fe favors the in-plane magnetic easy axis. Besides, its magnitude is significantly larger than that of the spin-conservation term. Therefore, the sum of the spin-conservation term and the spin-flip term is also negative with increasing $S$, as shown in Fig. 4.
Figure 4: Contribution to the MCA energy originating from the orbital moment anisotropy (spin-conservation term, blue) and the charge-density anisotropy (spin-flip term, red). The sum of the two terms (total MCA energy) is also shown (green).
@@
The present results also indicate that the PMA of $L1_{0}$ FePt thin films mainly originates from the Pt sites. Our previous XMCD study at the Pt $L_{2,3}$ edge Ikeda et al. (2017) shows that the OMA of Pt is negative ($m_{\text{orb}}^{\perp}<m_{\text{orb}}^{\parallel}$), suggesting that the OMA favors the in-plane magnetic easy axis. The study also suggests that $m_{\text{spin}}^{\text{eff}}$ of Pt along the out-of-plane direction is slightly larger than that along the in-plane direction for highly ordered samples, which might be explained by the increase of $m_{\text{T}}^{\perp}-m_{\text{T}}^{\parallel}$ with increasing $S$. Ikeda et al. (2017) According to Eq. (3), this has a positive contribution to PMA. Therefore, it follows that the PMA of the $L1_{0}$ FePt thin films is primarily due to the orbital asphericity of the Pt sites. These results are consistent with the theoretical calculation using the Green’s function method Solovyev, Dederichs, and Mertig (1995). To verify this hypothesis, AD-XMCD experiments at the Pt $L_{2,3}$ edge using hard x rays could provide a direct evidence, which remain to be performed in the future.
In summary, we have performed AD-XMCD measurements of FePt thin films with various $L1_{0}$ order degrees $S$. The MCA energy $K_{\text{u}}$ and the electric quadrupole moment $\braket{Q_{zz}}$, i.e. the charge-density anisotropy, have been quantitatively deduced from the angular dependence of XMCD. The deduced $\braket{Q_{zz}}$ is negative for the films with larger $S$, suggesting that the Fe 3dd orbitals elongated along the $z$ direction are preferentially occupied in the $L1_{0}$-ordered films. The MCA energy due to this charge-density anisotropy of Fe ions is negative and dominates over the positive contribution by OMA, consistent with the previous first-principles calculation Solovyev, Dederichs, and Mertig (1995). Considering that the OMA of Pt also favors the in-plane magnetic easy axis, Ikeda et al. (2017) one has to take into account the charge-density anisotropy of Pt in order to explain the large PMA of $L1_{0}$-ordered FePt thin films.